The perimeter of the rectangle is 138 inches. The length of the rectangle is 4x inches and the width is 3x+6 inches. Find the value of x
Answer by addingup(3677)

2L+2W = 138 Substitute for L and W using info provided in the problem:
2(4x)+2(3x+6) = 138
8x+6x+12 = 138
14x = 126
x = 9
:
Happy learning,
John